# CUORE Collaboration

### v1

@@ Abstract
We present the analysis techniques developed to explore the keV-scale energy region of the CUORE experiment, based on more than 2 tonne yr of data collected over 5 years. By prioritizing a stricter selection over a larger exposure, we are able to optimize data selection for thresholds at 10 keV and 3 keV with 691 kg yr and 11 kg yr of data, respectively. We study how the performance varies among the 988-detector array with different detector characteristics and data taking conditions. We achieve an average baseline resolution of 2.54 $\pm$ 0.14 keV FWHM and 1.18 $\pm$ 0.02 keV FWHM for the data selection at 10 keV and 3 keV, respectively. The analysis methods employed reduce the overall background by about an order of magnitude, reaching 2.06 $\pm$ 0.05 counts/(keV kg days) and 16 $\pm$ 2 counts/(keV kg days) at the thresholds of 10 keV and 3 keV. We evaluate for the first time the near-threshold reconstruction efficiencies of the CUORE experiment, and find these to be 26 $\pm$ 4 % and 50 $\pm$ 2 % at 3 keV and 10 keV, respectively. This analysis provides crucial insights into rare decay studies, new physics searches, and keV-scale background modeling with CUORE. We demonstrate that tonne-scale cryogenic calorimeters can operate across a wide energy range, from keV to MeV, establishing their scalability as versatile detectors for rare event and dark matter physics. These findings also inform the optimization of future large mass cryogenic calorimeters to enhance the sensitivity to low-energy phenomena.
@@ I Introduction
An optimization of CUORE’s data processing is necessary to make use of the keV-scale data. The threshold of CUORE detectors is of the order of a few keV, so the spectrum we are interested in is polluted by noise and spurious events unless additional event selection criteria are applied. To explore this energy region, we developed dedicated analysis methods devoted to the selection of the highest quality data close to the trigger thresholds. As an example, we show in Fig. 1 four events with a reconstructed energy below 150 keV. These pulses demonstrate how the noise contribution becomes more prominent at lower energies. This affects several steps of the analysis, primarily the estimation of the pulse shape parameters needed to characterize an event and the detector’s efficiencies. Moreover, this region of the energy spectrum is highly susceptible to spurious pulses, which originate from external sources, e.g. electronics and vibrations, inducing phonon excitation in the crystals.
@@
Beyond paving the way to new physics investigation at the keV-scale with CUORE, these results inform planning and goals for future stages of large mass cryogenic calorimeters. In particular, some relevant upgrades are foreseen to mitigate the vibration environment of the CUORE cryogenic facility in view of the next generation experiment CUPID (CUORE Upgrade with Particle IDentification) [18, 19]. Scintillating crystals, coupled to light detectors, will enable the dual readout of heat and light for particle identification in CUPID. Light detectors will be made of thin germanium cryogenic calorimeters. Such detectors are sensitive to vibration induced by the pulse tubes cryocoolers that are needed to cool down about a tonne of material down to 4K. The upgrades will set the stage for a second phase of CUORE with a low vibration-induced noise, ideal to focus the data taking for keV-scale studies. The physics potential of the second CUORE phase, as well as of CUPID, can rely on the results presented in this work.
This paper is organized as follows. Sec. II briefly describes the CUORE experiment and its data taking. Sec. III introduces the triggering procedure applied to data, a fundamental step for reconstructing keV-scale events. Sec. IV describes all the analysis strategies and techniques that we developed specifically for low energy deposits, focusing on the rejection of spurious events polluting the spectrum. Sec. V reports the results of the developed techniques in terms of selected exposure and energy resolution. Here we also investigates patterns in detector performance in the CUORE array over the course of data taking. Sec. VI details the verifications we performed about the energy reconstruction down to 3 keV. Sec. VII reports the estimate of the energy-dependent efficiency due to event reconstruction and selection. Finally, Sec. VIII shows the energy spectrum resulting from this study and highlights its main features.
@@ II The CUORE Experiment
CUORE, located in the Laboratori Nazionali del Gran Sasso (LNGS) in Assergi, Italy, is composed of 988 5 cm $\times$ 5 cm $\times$ 5 cm $\mathrm{natural}$ $\mathrm{TeO_{2}}$ crystal cryogenic calorimeters. The crystals provide 742 kg of active mass and they are arranged in an array of 19 towers with 13 floors each mounted by copper holders secured by polytetraflouroethylene (PTFE). Towers are placed such that 12 of them form an outer layer containing the remaining 7 ones.
@@
III Optimum Trigger
The first CUORE data processing step is to apply an offline trigger, the Optimum Trigger (OT) [27, 9], to the continuous data-stream to improve the energy threshold. In particular, it triggers the waveforms upon optimizing the signal-to-noise ratio through digital filtering. For this purpose, a matched filter, the Optimum Filter (OF) [28, 29], is used to suppress the frequencies most affected by the noise. This algorithm employs a detector response based on the average signal pulse shape and the average detector noise power spectrum extracted from data.
Figure 2: Distribution of the trigger thresholds for the detector-dataset pairs used in the 2 tonne yr low energy dataset. The different colors highlight the pairs belonging to different data taking conditions, i.e. operation temperature at 12 or 15 mK, and employing an optimal vibration damping system configuration.
IV Data Analysis
Most of the data processing adopted for this study relies on the standard CUORE analysis chain, which is described in Ref. [30]. Before being triggered, a waveform’s noise is de-correlated [31] from auxiliary devices (microphones, low frequency accelerometers and seismometers) placed in the cryostat hut, in order to mitigate the environmental contribution to vibrational noise [31]. Once triggered, waveforms are processed with the OF, as described in Sec. III, but with an improved evaluation of the template of the detector response. The amplitude estimate is corrected by the thermal gain to account for temperature instabilities. Finally, detectors are energy-calibrated by using dedicated runs with 232Th - 60Co source.
@@ IV Data Analysis
IV.1 Tagging of multi-site events
@@
IV.2 Spurious events discrimination
@@
For this purpose, we define two energy intervals, the Region Of Interest (ROI) and the Region Of Reference (ROR): the first determines the lowest energy threshold we aim to analyze, while the second is a higher energy interval to be used as a performance comparison. While ROI and ROR are free variables of the algorithm, their specific values for this analysis will be defined in Sec. V.
@@
Figure 4: Example of $\chi^{2}_{OF}$ (Pulse Shape Parameter) as a function of the energy for a single detector and dataset. We show two pulses having similar energy ($\sim$15 keV), but different $\chi^{2}_{OF}$. The left pulse belongs to the spurious events band, while the right one to the signal band.
@@
We scan along several cut levels on the $\chi^{2}_{OF}$ variable and we apply it to all the detectors. We define the cut levels by dividing the $\chi^{2}_{OF}$ distribution of each detector in equally-spaced quantiles, using data from the 511 keV peak. The cut level is the actual quantile of the $\chi^{2}_{OF}$ distribution. In this way, while sharing the same cut level, each detector has a different cut on the absolute $\chi^{2}_{OF}$ value, which is dependent on its distribution.
Then, we evaluate the signal efficiency $\varepsilon_{PS}$ by fitting the Te X-ray events selected and rejected by the cut simultaneously. We also count the number of events left by the cut in the ROI ($N_{B}$), assuming they are dominated by physical background or spurious pulses.
We identify the optimal cut as the one that maximizes the ratio $\varepsilon_{PS}/N_{B}$, in order to balance the loss in efficiency with the gain in background level consequent to a stronger cut.
@@
The pulse shape reconstruction by the $\chi^{2}_{OF}$ variable is inefficient when the energy deposit is close to the noise threshold of the detector. Thus, it is not possible to distinguish signals from spurious pulses below a given energy, which depends on the detector. We report a few examples in Fig. 5. Once the $\chi^{2}_{OF}$ cut is applied, we can detect the presence of unwanted events by looking at two features:
  • Event Clustering: when the spurious and signal bands merge in the $\chi^{2}_{OF}$ vs Energy plot, the number of events rises at lower energies. We quantify this effect by building a variable called Purity ($\mathcal{P}$), namely the fraction of signal events in the ROI. Assuming a similar density of physical events at different energies, we estimate $\mathcal{P}$ as $N_{S}/N_{B}$, where $N_{S}$ is the number of events in the ROR scaled by the energy width of the ROI. If the ROR is well chosen, after the $\chi^{2}_{OF}$ cut it contains a pure sample of physics events. The assumption of having the same physical events density at different energies does not affect the detector selection, since we aim to classify and order them from the most to the least contaminated by spurious events.
@@
In the following, we adopt a ROR between 30 and 50 keV. This choice is led by the necessity to maintain consistent pulse shape characteristics and event rates with respect to the ROI. The second requirement is no longer satisfied above $\sim$50 keV due to an observed difference by about an order of magnitude of the background level, as better outlined in Sec. VIII.
@@
V Data Selection Results
We apply the previously-described methods on the 2 tonne yr CUORE data release using two different ROIs: [10,20] keV, of interest for solar axions searches [36, 10, 6, 7], and [3,10] keV. The latter is the lowest accessible energy region of CUORE since no detectors achieved a trigger threshold lower than 2 keV. Moreover, it is of interest to study the excess of events at 4.7 keV already observed in past CUORE demonstrators [37].
We achieve a total of 11 and 691 kg yr TeO2 exposure for detectors available down to 3 keV and 10 keV, respectively. This amounts to 0.5% and 34% of the total collected exposure, respectively. Considering only 12 mK datasets, the exposure at 3 keV reaches 1.2% of the total one.
@@ V Data Selection Results
The selected exposure, or number of detector-run pairs, by dataset has various trends when compared to data taking conditions, as mentioned in Sec. I, such as detector temperature and vibration damping configuration, as shown in Fig. 9. The operation temperature heavily affects the exposures available at 3 keV, as no detectors pass selections at this energy in datasets operated at a warmer 15 mK temperature. The time periods labeled as “damping” in Fig. 9 and 10 correspond to an optimal configuration of the custom vibration isolators made by Minus K Technology [21]. This system is located under the Y-beam that supports the CUORE Top Support Plate of the detector and it is meant to mitigate sub-Hz vibration noise, with a cut-off at 0.5 Hz[20]. In particular, we find that a specific configuration operated over a few datasets is beneficial to mitigate the noise at very low energy, leading in a broader subset of detectors usable down to 3 keV.
The same trend along the three phases of data taking described in Sec. I holds when considering energy resolution, as it improves with colder temperatures and the optimal damping configuration. We compute the baseline resolution by fitting to a Gaussian function the energy distribution of randomly triggered noise events with no pulses in the window, and getting the FWHM resolution.
@@
VI Calibration validation at low energy
VII Efficiencies
On each dataset, we evaluate efficiencies for anti-coincidence and $\chi^{2}_{OF}$ cuts, and for reconstruction on the detectors selected at 10 and 3 keV.
@@ VII Efficiencies
In previous CUORE analyses, only pulsers with an average energy of $\sim$3 MeV, were used to calculate these efficiencies [34, 1, 4]. We adapt the same algorithm to run on lower energy pulser events to evaluate the near-threshold reconstruction efficiencies and evaluate their energy dependence. In particular, we evaluate the reconstruction efficiencies on a dataset basis, for every group of pulser events at decreasing energy from $\sim$350 keV down to $\sim$5 keV. Then, for each dataset, we fit each of the reconstruction efficiencies as a function of the energy to a logistic function. The resulting curve for each efficiency is averaged among all the datasets weighted by exposure. Finally, the overall reconstruction efficiency is produced by analytically multiplying all the exposure-weighted average efficiencies. The result is shown for the detectors selected at 3 keV and 10 keV in Fig. 12.
@@
We assume the 40K line is a pure M1 event, and we consider the presence of the previously described $\sim$3.2 keV X-ray to be negligible since its emission intensity is as low as $\sim$0.07[33]. Without knowledge of known physical structures in the spectrum to evaluate these efficiencies, we assume them to be constant along the energy scale of interest for this work. The anti-coincidence efficiency is found to be compatible with 1 in all cases and so it doesn’t affect estimations of the overall efficiency.
VIII Low Energy Spectra
In Fig. 13 and 14 we show the exposure-weighted low energy background spectrum of the 2 tonne yr of CUORE data taking down to 10 and 3 keV, respectively. The initial cuts require the trigger threshold to be less than the chosen lower edge, either 3 or 10 keV (spectrum in yellow). Next, we apply the detector-dependent cut on $\chi^{2}_{OF}$ (in blue), and lastly the detector selections (in green). The blue and green spectra are weighted by the $\chi^{2}_{OF}$ cut efficiency. In this way, it is possible to compare the background level reduction due to the rejection of spurious pulses before and after the cut. The $\chi^{2}_{OF}$ cut and the detector selection show the greatest difference from basic cuts below 30 keV, reducing the background by an order of magnitude. This region is the most interesting for multiple planned and potential physics searches. The result of this work highlights the main spectral features of CUORE at low energy.

### v2

@@ Abstract
We present the analysis techniques developed to explore the keV-scale energy region of the CUORE experiment, based on more than 2 tonne yr of data collected over 5 years. By prioritizing a stricter selection over a larger exposure, we are able to optimize data selection for thresholds at 10 keV and 3 keV with 691 kg yr and 11 kg yr of data, respectively. We study how the performance varies among the 988-detector array with different detector characteristics and data taking conditions. We achieve an average baseline resolution of 2.54 $\pm$ 0.14 keV FWHM and 1.18 $\pm$ 0.02 keV FWHM for the data selection at 10 keV and 3 keV, respectively. The analysis methods employed reduce the overall background by about an order of magnitude, reaching 2.06 $\pm$ 0.05 counts/(keV kg days) and 16 $\pm$ 2 counts/(keV kg days) at the thresholds of 10 keV and 3 keV. We evaluate for the first time the near-threshold reconstruction efficiencies of the CUORE experiment, and find these to be 50 $\pm$ 2 % and 26 $\pm$ 4 % at 10 keV and 3 keV, respectively. This analysis provides crucial insights into rare decay studies, new physics searches, and keV-scale background modeling with CUORE. We demonstrate that tonne-scale cryogenic calorimeters can operate across a wide energy range, from keV to MeV, establishing their scalability as versatile detectors for rare event and dark matter physics. These findings also inform the optimization of future large mass cryogenic calorimeters to enhance the sensitivity to low-energy phenomena.
@@ I Introduction
[…]
@@
This paper is organized as follows. Sec. II briefly describes the CUORE experiment and its data taking. Sec. III describes all the analysis strategies and techniques that we developed specifically for low energy deposits. In particular, Sec. III.1 details the verifications we performed about the energy reconstruction down to 3 keV, Sec. III.2 introduces the offline triggering procedure applied to data, Sec. III.3 highlights the improvements on tagging multi-sites events at the keV-scale, and Sec III.4 focuses on the rejection of spurious events polluting the near-threshold spectrum. Sec. IV reports the results of the developed techniques in terms of selected exposure and energy resolution. Here we also investigates patterns in detector performance in the CUORE array over the course of data taking. Sec. V reports the estimate of the energy-dependent efficiency due to event reconstruction and selection. Finally, Sec. VI shows the energy spectrum resulting from this study and highlights its main features.
@@ II The CUORE Experiment
CUORE, located in the Laboratori Nazionali del Gran Sasso (LNGS) in Assergi, Italy, is composed of 988 5 cm $\times$ 5 cm $\times$ 5 cm $\mathrm{natural}$ $\mathrm{TeO_{2}}$ crystal cryogenic calorimeters. The crystals provide 742 kg of active mass and they are arranged in an array of 19 towers with 13 floors each mounted by copper holders secured by polytetraflouroethylene (PTFE). Towers are placed such that 12 of them form an outer layer containing the remaining 7 ones [47].
@@
III Data Analysis
Most of the data processing adopted for this study relies on the standard CUORE analysis chain, which is described in Ref. [49]. Before being triggered, a waveform’s noise is de-correlated [50] from auxiliary devices (microphones, low frequency accelerometers and seismometers) placed in the cryostat hut, in order to mitigate the environmental contribution to vibrational noise [50]. Once triggered, waveforms are processed with the Optimum Filter (OF) [34, 44], a matched filter used to suppress the frequencies most affected by the noise. This algorithm employs a detector response based on the average signal pulse shape and the average detector noise power spectrum extracted from data to build the transfer function on a detector basis. The amplitude estimate is corrected by the thermal gain to account for temperature instabilities. Finally, detectors are energy-calibrated by using dedicated runs with 232Th - 60Co source.
@@ III.1 Validation data and energy scale calibration
III.2 Optimum Trigger
The first CUORE data processing step is to apply an offline trigger, the Optimum Trigger (OT) [32, 8], to the continuous data-stream to identify events with an improved energy threshold. In particular, it triggers the waveforms upon optimizing the signal-to-noise ratio through digital filtering with the OF. The filter transfer function is built for each CUORE detector from derivative triggered data and applied on its continuous data stream; subsequently events are triggered. A signal is flagged by the OT if the filtered amplitude exceeds the filtered noise resolution by a factor $n$. The parameter $n$ is configurable on a detector basis, while the filtered noise resolution is known a priori from the filter transfer function of the given detector [26].
Figure 4: Distribution of the trigger thresholds namely, the energy at which trigger efficiency is 90%, for the detector-dataset pairs used in the 2 tonne yr low energy dataset. The different colors highlight the pairs belonging to different data taking conditions, i.e. operation temperature at 12 or 15 mK, and employing an optimal vibration damping system configuration. Any distribution refers to all the operated CUORE detectors, but to a different number of datasets, depending on the experimental conditions.
III.3 Tagging of multi-site events
@@
III.4 Spurious events discrimination
@@
For this purpose, we define two energy intervals, the Region Of Interest (ROI) and the Region Of Reference (ROR): the first determines the lowest energy threshold we aim to analyze, depending on the purpose of the study, while the second is a higher energy interval to be used as a performance comparison. While ROI and ROR are free variables of the algorithm, their specific values for this analysis will be defined in Sec. IV.
@@
Figure 5: Example of $\chi^{2}_{OF}$ (Pulse Shape Parameter) as a function of the energy for a single detector, with a $\sim$2 keV trigger threshold, and dataset. We show two pulses having similar energy ($\sim$15 keV), but different $\chi^{2}_{OF}$. The left pulse belongs to the spurious events band, while the right one to the signal band.
@@
Following this consideration, we apply a pulse shape cut to select events with $\chi^{2}_{OF}$ lower than a given value. We scan along several cut values to identify the optimal one. For each value we estimate the signal efficiency ($\varepsilon_{PS}$), by fitting the Te X-ray events selected and rejected by the cut simultaneously, and we count the number of events left by the cut in the ROI ($N_{B}$), assuming they are dominated by physical background or spurious pulses. We identify the optimal cut as the one that maximizes the ratio $\varepsilon_{PS}/N_{B}$, in order to balance the loss in efficiency with the gain in background level consequent to a stronger cut.
To account for different performance along the CUORE array, we optimize the pulse shape cut independently for each detector. For this purpose, for each detector we define the cut values by looking at the $\chi^{2}_{OF}$ distribution of a given physical peak (namely, the 511 keV, which is a compromise between low energy and good signal-to-noise ratio). In particular, we divide it in intervals of equal probability (quantiles): the cut values on which we perform the scan are the $\chi^{2}_{OF}$ values corresponding to the edges of the quantiles, which range from 20 to 90 % with a 5% steps. This allows to scan over $\chi^{2}_{OF}$ values with different scales on different detectors.
@@
The pulse shape reconstruction by the $\chi^{2}_{OF}$ variable is inefficient when the energy deposit is close to the noise threshold of the detector. Thus, it is not possible to distinguish signals from spurious pulses below a given energy, which depends on the detector. We report a few examples in Fig. 6. Once the $\chi^{2}_{OF}$ cut is applied, we can identify detectors which still present unwanted events by looking at two features:
  • Event Clustering: when the spurious and signal bands merge in the $\chi^{2}_{OF}$ vs Energy plot, the number of events rises at lower energies. We quantify this effect by building a variable called Purity ($\mathcal{P}$), namely the fraction of signal events in the ROI. Assuming a similar density of physical events at different energies, we estimate $\mathcal{P}$ as $N_{S}/N_{B}$, where $N_{S}$ is the number of events in the ROR scaled to the energy width of the ROI. If the ROR is well chosen, after the $\chi^{2}_{OF}$ cut it contains a pure sample of physics events. The assumption of having the same physical events density at different energies does not affect the detector selection, since we aim to classify and order them from the most to the least contaminated by spurious events.
@@
In the following, we adopt a ROR between 30 and 50 keV. This choice is led by the necessity to maintain consistent pulse shape characteristics and event rates with respect to the ROI. The second requirement is no longer satisfied above $\sim$50 keV due to an observed difference by about an order of magnitude of the background level, as better outlined in Sec. VI.
@@
IV Data Selection Results
We apply the previously-described methods on the 2 tonne yr CUORE data release using two different ROIs: [10,20] keV, of interest for solar axions searches [46, 35, 36, 37], and [3,10] keV. The latter is the lowest accessible energy region of CUORE since no detectors achieved a trigger threshold lower than 2 keV. Moreover, it is of interest to study the excess of events at 4.7 keV already observed in past CUORE demonstrators [13]. The $\chi^{2}_{OF}$ cut optimization and the detectors selection is repeated for each chosen ROIs. We achieve a total of 691 and 11 kg yr TeO2 exposure for detectors available down to 10 keV and 3 keV, respectively. This amounts to 34% and 0.5% of the total collected exposure, respectively. Considering only 12 mK datasets, the exposure at 3 keV reaches 1.2% of the total one.
@@ IV Data Selection Results
The selected exposure, or number of detector-run pairs, by dataset has various trends when compared to data taking conditions, as mentioned in Sec. I, such as detector temperature and vibration damping configuration, as shown in Fig. 10. The operation temperature heavily affects the exposures available at 3 keV, as no detectors pass selections at this energy in datasets operated at a warmer 15 mK temperature. The time periods labeled as “damping” in Fig. 10 and 11 correspond to an optimal configuration of the custom vibration isolators made by Minus K Technology [38]. This system is located under the Y-beam that supports the CUORE Top Support Plate of the detector and it is meant to mitigate sub-Hz vibration noise, with a cut-off at 0.5 Hz[9]. In particular, we find that a specific configuration operated over a few datasets is beneficial to mitigate the noise at very low energy, leading in a broader subset of detectors usable down to 3 keV.
The same trend along the three phases of data taking described in Sec. I holds when considering energy resolution, as it improves with colder temperatures and the optimal damping configuration. We compute the baseline resolution by fitting to a Gaussian function the energy distribution of randomly triggered noise events with no pulses in the window, and getting the FWHM resolution. In the optimal data taking configuration we can appreciate the larger Te X-rays peak width due to its composition made of multiple lines.
@@
V Efficiencies
On each dataset, we evaluate efficiencies for anti-coincidence and pulse shape (PS) cuts, and for reconstruction on the detectors selected at 10 and 3 keV.
@@ V Efficiencies
In previous CUORE analyses, only pulsers with an average energy of $\sim$3 MeV, were used to calculate these efficiencies [1, 47, 48]. We adapt the same algorithm to run on lower energy pulser events to evaluate the near-threshold reconstruction efficiencies and evaluate their energy dependence. In particular, we evaluate the reconstruction efficiencies on a dataset basis, for every group of pulser events at decreasing energy from $\sim$350 keV down to $\sim$5 keV. Then, for each dataset, we fit each of the reconstruction efficiencies as a function of the energy to a logistic function. The resulting curve for each efficiency is averaged among all the datasets weighted by exposure. Finally, the overall reconstruction efficiency is produced by analytically multiplying all the exposure-weighted average efficiencies. The result is shown for the detectors selected at 10 keV and 3 keV in Fig. 12.
@@
We assume the 40K line is a pure M1 event, and we consider the presence of the previously described $\sim$3.2 keV X-ray to be negligible since its emission intensity is as low as $\sim$0.07% [28]. Without knowledge of known physical structures in the spectrum to evaluate these efficiencies, we assume them to be constant along the energy scale of interest for this work. The anti-coincidence efficiency is found to be compatible with 1 in all cases and so it doesn’t affect estimations of the overall efficiency.
VI Low Energy Spectra
In Fig. 13 and 14 we show the exposure-weighted low energy background spectrum of the 2 tonne yr of CUORE data taking down to 10 and 3 keV, respectively.
The initial cuts require the trigger threshold to be less than the chosen lower edge, either 10 or 3 keV (spectrum in yellow). Next, we apply the detector-dependent cut on $\chi^{2}_{OF}$ (in blue), and lastly the detector selections (in green). The blue and green spectra are weighted by the $\chi^{2}_{OF}$ cut efficiency. In this way, it is possible to compare the background level reduction due to the rejection of spurious pulses before and after the cut. The $\chi^{2}_{OF}$ cut and the detector selection show the greatest difference from basic cuts below 30 keV, reducing the background by an order of magnitude. This region is the most interesting for multiple planned and potential physics searches.
The result of this work highlights the main spectral features of CUORE at low energy which are summarized in Table 1.
Energy [keV] Multiplicity Origin
$\sim$ 27, 31 M2 Te X-rays
$\sim$46 M1 210Pb $\beta+\gamma$
$\sim$90 M1, M2 210Po nuclear recoil
$\sim$145 M1 125mTe
Energy [keV] Multiplicity Rate [counts/(kg d)]
$\sim$4.7 M1 19.1 $\pm$ 1.0
$\sim$10 M1 0.66 $\pm$ 0.07
$\sim$13 M1 0.67 $\pm$ 0.07
$\sim$31 M1 1.35 $\pm$ 0.04
$\sim$36 M1 8.3 $\pm$ 0.3
Table 1: Summary of the main features of CUORE energy spectrum up to 200 keV. The top table reports structures of known origin, while the bottom one those which are currently under further investigation.